If a man 20 kg move on plank of 60 kg find displacement of plank and man.
1
Expert's answer
2020-10-23T11:55:29-0400

Explanations & Calculations





  • To make sense this system should lie on a frictionless surface for the plank to move on.
  • All the motions take place here are internal, hence the status of the center of mass (CM) is unchanged.
  • Any movement occur in such a way that the initial position of the CM is kept fixed in place.
  • Consider Earth's frame of reference for measuring distances.
  • Get the distance "\times" mass product,

"\qquad\qquad
\begin{aligned}
\small (20+60)\times d_{CM} &= \small 20\times 0+60\times \frac{L}{2}\\
\small d_{CM} & =\small \frac{3L}{8}
\end{aligned}"

  • If the man walks some x distance towards the other end & at the same time the planks moves some y distance in the backward direction then it follows the concept stated above. Therefore,

"\qquad\qquad
\begin{aligned}
\small 80\times \frac{3L}{8}&= \small 20\times x+60\times (\frac{l}{2}-y)\\
\small \bold{ y} &= \bold{\small \frac{x}{3}} 
\end{aligned}"

  • By the input of relevant measurements, distances travelled by each could be found.


Example:

  • If the man reaches the other end then the distance he travelled becomes "\small x = L-y"
  • Then the plank travels,

"\qquad\qquad
\begin{aligned}
\small y &= \small \frac{L-y}{3}\\
\small y&= \small \bold{\frac{L}{4}}
\end{aligned}"